(10 pts) A car starts on a trip and travels at a speed of 55 mph. Two hours later, a second car starts on the same trip and travels at a speed of 65 mph.
When the second car has been on the road for t hours, the first car has traveled ____ miles and the second car has traveled ____miles.
At time t the distance between the first car and the second car is_____ miles.
The ratio of the distance the second car has traveled and the distance the first car has traveled is ______.
The second car catches up with the first car _____ hours after the departure of the first car. (Those are some determined drivers!)

Answer by Paul(988)

(10 pts) A car starts on a trip and travels at a speed of 55 mph. Two hours later, a second car starts on the same trip and travels at a speed of 65 mph.
let t be the time
55(t+2)=65t
55t+110=65t
110=10t
t=11
11+2 = 13